# Accelerating Eigenvalue Dataset Generation via Chebyshev Subspace Filter

### v1

@@ C.1 Overview
The general methodology for solving the eigenvalue problems of differential operators numerically, employing techniques such as Finite Difference Method (FDM), Finite Element Method (FEM), and Spectral Method, can be delineated through the following pivotal steps (Strikwerda, 2004; Hughes, 2012; Johnson, 2012; LeVeque, 2002):
@@ D.1 Baseline
  • Jacobi-Davidson (JD): Implemented in SLEPc (v3.21.1). The implementation uses ’bcgsl’ as the linear equation solver, ’bjacobi’ as the preconditioner, and sets the linear equation solving precision to 1e-5.
@@ 1. Generalized Poisson Operator
We consider two-dimensional generalized Poisson operators, which can be described by the following equation (Li et al., 2020; Rahman et al., 2022; Kovachki et al., 2021; Lu et al., 2022):
@@ E.1 Main Experimental Data
Table 8: Comparison of average computation times (in seconds) for eigenvalue problems using various algorithms on the Helmholtz operator dataset. The first row lists different algorithms, and the first column shows the number of eigenvalues $L$ computed for each matrix. Matrix dimension = 6400, precision = 1e-8. The symbol ’-’ denotes data not recorded due to excessive computation times.
@@
Table 9: Comparison of average computation times (in seconds) for eigenvalue problems using various algorithms on the Vibration operator dataset. The first row lists different algorithms, and the first column shows the number of eigenvalues $L$ computed for each matrix. Matrix dimension = 10000, precision = 1e-8. The symbol ’-’ denotes data not recorded due to excessive computation times.
@@ E.2 Analysis of the Influence of Matrix Dimension
Table 10: Comparison of different algorithms’ computation time (in seconds) for varying matrix dimensions using the generalized Poisson operator dataset. Results show average computation times for solving 400 eigenvalues with a precision of 1e-12.
@@ E.3 Analysis of Computational Times for SCSF Components
We conducted a statistical analysis of the average time consumption for each component of the SCSF algorithm on the generalized Poisson operator dataset, with a matrix dimension of 2500 and the number of eigenvalues to be solved set to 100. The results are presented in Table 11. The notation "line x" within parentheses corresponds to line x in Algorithm 3, "ALL" denotes the total time consumption, and "sort" represents the average time required by the sorting algorithm. It is evident that the filtering process accounts for over 70% of the total time consumption, which aligns with our theoretical analysis in Section 3.2.
@@ E.5 Reliability of Generated Data as Ground Truth
A key concern was whether the data generated by our method, which relies on numerical solvers, is a reliable "ground truth" for training neural networks. To address this, we trained a NeurKItt (Luo et al., 2024) model on generalized Poisson datasets generated by various solvers (including our SCSF) at different matrix dimensions. The precision for all solvers was set to a high tolerance of $10^{-12}$.
@@ E.6 Comparison with Supervised and Unsupervised Neural Network Methods
To clarify the significance of accelerating dataset generation for the dominant supervised learning paradigm, we conducted an experiment comparing the performance and resource trade-offs of different categories of eigensolvers. We evaluated our method (SCSF), a traditional solver (Eigsh), a supervised neural network (NeurKItt (Luo et al., 2024)), and two state-of-the-art unsupervised neural networks (NeuralEF (Deng et al., 2022) and NeuralSVD Ryu et al. (2024)) on a 2D Helmholtz problem (solving for the smallest 100 eigenvalues, matrix dimension $6400$).
Table 16: Comparison of different eigensolver paradigms on a 2D Helmholtz problem. "Solving Time" for unsupervised methods refers to the entire optimization process required to find the solution for a single operator instance.
@@
The results, presented in Table 16, highlight the distinct characteristics of each approach. Unsupervised methods obviate the need for pre-generated datasets but require substantial "solving time" for each new operator, as they essentially perform an optimization from scratch. In contrast, supervised methods offer near-instantaneous inference but demand significant upfront investment in both data generation and model training. Our method, SCSF, dramatically reduces the data generation bottleneck for these powerful supervised models.
@@ E.8 Analysis of Failure Cases: Discontinuous Datasets
The core assumption of SCSF is that the dataset is generated from a process with underlying continuity, allowing our sorting algorithm to group similar problems effectively. To investigate the behavior of SCSF when this assumption is violated, we simulated an extreme failure case by creating a discontinuous dataset. We mixed problems from the Helmholtz and Poisson datasets (1:1 ratio, dimension $6400$, $L=200$ eigenvalues) and solved them sequentially.
@@
Metric Eigsh LOBPCG ChFSI SCSF (w/o sort) SCSF
Time (s) 154 280 132 118 98
Table 18 presents the results. As expected, the performance gain of SCSF is reduced in this discontinuous scenario because the inter-problem correlation is disrupted, diminishing the effectiveness of the sorting module. However, even in this challenging case, SCSF still provides a notable speedup over baseline solvers, demonstrating a degree of robustness. The performance of ‘SCSF‘ approaches that of ‘SCSF (w/o sort)‘, confirming that the sorting component’s benefit is tied to dataset continuity.

### v2

@@ C.1 Overview
The general methodology for solving the eigenvalue problems of differential operators numerically, employing techniques such as Finite Difference Method (FDM), Finite Element Method (FEM), and Spectral Method, can be delineated through the following pivotal steps (Strikwerda, 2004; Hughes, 2012; Johnson, 2012; LeVeque, 2002; Wei et al., 2024):
@@ D.1 Baseline
  • Jacobi-Davidson (JD): Implemented in SLEPc (v3.21.1). The implementation uses ‘bcgsl’ as the linear equation solver, ‘bjacobi’ as the preconditioner, and sets the linear equation solving precision to 1e-5.
@@ 1. Generalized Poisson Operator
We consider two-dimensional generalized Poisson operators, which can be described by the following equation (Li et al., 2020; Rahman et al., 2022; Kovachki et al., 2021; Lu et al., 2022; Gao et al., 2025):
@@ E.1 Main Experimental Data
Table 8: Comparison of average computation times (in seconds) for eigenvalue problems using various algorithms on the Helmholtz operator dataset. The first row lists different algorithms, and the first column shows the number of eigenvalues $L$ computed for each matrix. Matrix dimension = 6400, precision = 1e-8. The symbol ‘-’ denotes data not recorded due to excessive computation times.
@@
Table 9: Comparison of average computation times (in seconds) for eigenvalue problems using various algorithms on the Vibration operator dataset. The first row lists different algorithms, and the first column shows the number of eigenvalues $L$ computed for each matrix. Matrix dimension = 10000, precision = 1e-8. The symbol ‘-’ denotes data not recorded due to excessive computation times.
@@ E.2 Analysis of the Influence of Matrix Dimension
Table 10: Comparison of different algorithm computation time (in seconds) for varying matrix dimensions using the generalized Poisson operator dataset. Results show average computation times for solving 400 eigenvalues with a precision of 1e-12.
@@ E.3 Analysis of Computational Times for SCSF Components
We conducted a statistical analysis of the average time consumption for each component of the SCSF algorithm on the generalized Poisson operator dataset, with a matrix dimension of 2500 and the number of eigenvalues to be solved set to 100. The results are presented in Table 11. The notation ‘line x’ within parentheses corresponds to line x in Algorithm 3, ‘ALL’ denotes the total time consumption, and ‘sort’ represents the average time required by the sorting algorithm. It is evident that the filtering process accounts for over 70% of the total time consumption, which aligns with our theoretical analysis in Section 3.2.
@@ E.5 Reliability of Generated Data as Ground Truth
A key concern was whether the data generated by our method, which relies on numerical solvers, is a reliable ‘ground truth’ for training neural networks. To address this, we trained a NeurKItt (Luo et al., 2024) model on generalized Poisson datasets generated by various solvers (including our SCSF) at different matrix dimensions. The precision for all solvers was set to a high tolerance of $10^{-12}$.
@@ E.6 Comparison with Supervised and Unsupervised Neural Network Methods
To clarify the significance of accelerating dataset generation for the dominant supervised learning paradigm, we conducted an experiment comparing the performance and resource trade-offs of different categories of eigensolvers. We evaluated our method (SCSF), a traditional solver (Eigsh), a supervised neural network (NeurKItt Luo et al. (2024)), and two state-of-the-art unsupervised neural networks (NeuralEF Deng et al. (2022) and NeuralSVD Ryu et al. (2024)) on a 2D Helmholtz problem (solving for the smallest 100 eigenvalues, matrix dimension $6400$).
Table 16: Comparison of different eigensolver paradigms on a 2D Helmholtz problem. ‘Solving Time’ for unsupervised methods refers to the entire optimization process required to find the solution for a single operator instance.
@@
The results, presented in Table 16, highlight the distinct characteristics of each approach. Unsupervised methods obviate the need for pre-generated datasets but require substantial ‘solving time’ for each new operator, as they essentially perform an optimization from scratch. In contrast, supervised methods offer near-instantaneous inference but demand significant upfront investment in both data generation and model training. Our method, SCSF, dramatically reduces the data generation bottleneck for these powerful supervised models.
@@ E.8 Analysis of Failure Cases: Discontinuous Datasets
The core assumption of SCSF is that the dataset is generated from a process with underlying continuity, allowing our sorting algorithm to group similar problems effectively. To investigate the behavior of SCSF when this assumption is violated, We simulated a gradual mixing of the two datasets (Helmholtz and Poisson). We varied the proportion of Helmholtz operators from 100% (a fully continuous dataset) down to 0% (a different fully continuous dataset), with mixed ratios in between representing varying degrees of discontinuity. (dimension $6400$, $L=200$ eigenvalues) and solved them sequentially.
@@
Helmholtz % Eigsh LOBPCG ChFSI SCSF (w/o sort) SCSF
100% (Helmholtz) 152 130 107 46 31
75% 154 203 116 94 78
50% (1:1 mix) 154 280 132 118 98
25% 152 359 141 108 80
0% (Poisson) 149 454 153 81 52
Table 18 presents the results. As expected, the performance gain of SCSF is reduced in this discontinuous scenario because the inter-problem correlation is disrupted, diminishing the effectiveness of the sorting module. However, even in this challenging case, SCSF still provides a notable speedup over baseline solvers, demonstrating a degree of robustness. The performance of ‘SCSF’ approaches that of ‘SCSF (w/o sort)’, confirming that the sorting component’s benefit is tied to dataset continuity.
@@ E.9 Cost-Benefit Analysis of the Sorting Algorithm
E.10 Sensitivity to Parameterization
To assess the sensitivity of the proposed parameter sorting to different PDE parameterizations, we conducted an additional experiment comparing two distinct discretization methods for the Helmholtz problem: the Finite Difference Method (FDM) (utilized in the main text) and the Finite Element Method (FEM/Galerkin).
As shown in Table 19, SCSF demonstrates consistent speedups across both discretization schemes. For the FEM dataset, we utilized a dimension of 10,000 with a tolerance of 1e-8.
Table 19: Comparison of SCSF performance under different parameterizations (Helmholtz). Times are in seconds. ‘-’ indicates non-convergence or excessive time cost.
Dataset $L$ Eigsh LOBPCG KS JD ChFSI SCSF (ours)
FDM (Central Diff) 200 151.7 129.9 98.34 489.6 107.1 31.31
400 253.5 460.4 283.0 3829 121.5 40.52
Dim: 6400, Tol: 1e-8 600 398.8 1031 329.6 - 146.2 51.32
FEM (Galerkin) 200 623.1 515.7 365.6 1885 450.9 116.7
400 938.5 1746 1166 - 481.3 142.9
Dim: 10000, Tol: 1e-8 600 1499 - 1299 - 599.9 196.3
These results confirm that SCSF remains highly effective regardless of the discretization scheme, provided the relationship between parameters and operators remains continuous.
E.11 Generalization of Truncation Level ($p_{0}$)
To justify this universal constant, we calculated the ratio of the Frobenius norm of high-frequency components (frequency >20>20) to the total norm for each dataset. The results are presented in Table 20.
Table 20: Ratio of high-frequency component energy (frequency $p_{0}=20$) to total energy across different PDE datasets.
Dataset Poisson Ellipse Helmholtz Vibration
High-frequency Ratio ($p_{0}$) 3.4% 3.7% 4.4% 4.8%
As shown, high-frequency components account for less than 5% of the total energy across all diverse PDE families. This confirms that $p_{0}=20$ consistently captures over 95% of the structural information inherent in the parameter matrices.
Appendix F Theoretical Analysis of Truncated FFT Sorting
Our sorting algorithm approximates the distance between operators to optimize the solving sequence. To analyze the theoretical validity of this approximation, we consider the isometric property of the Discrete Fourier Transform (DFT). Let $P^{(i)}$ and $P^{(j)}$ denote the discretized parameter matrices (e.g., diffusion coefficients) for two distinct problems. By Parseval’s identity, the DFT is a unitary transformation (up to a scaling factor), implying that the Frobenius distance in the spatial domain is strictly equivalent to that in the frequency domain:
$\|P^{(i)}-P^{(j)}\|_{F}^{2}=\|\text{FFT}(P^{(i)})-\text{FFT}(P^{(j)})\|_{F}^{2},$
where $\text{FFT}(\cdot)$ denotes the normalized DFT. The SCSF algorithm computes a proxy distance using a truncated spectrum, retaining only frequencies within a low-frequency bandwidth $p_{0}$. Consequently, the exact distance can be decomposed into the sorting metric and a residual truncation error:
$\|P^{(i)}-P^{(j)}\|_{F}^{2}=\underbrace{\|\text{Trunc}_{p_{0}}(\Delta\hat{P})\|_{F}^{2}}_{\text{SCSF Metric}}+\underbrace{\|(I-\text{Trunc}_{p_{0}})(\Delta\hat{P})\|_{F}^{2}}_{\text{Truncation Error }\epsilon_{ij}},$
where $\Delta\hat{P}=\text{FFT}(P^{(i)}-P^{(j)})$ and $\text{Trunc}_{p_{0}}$ is the truncation operator.
The magnitude of the error term $\epsilon_{ij}$ is governed by the spectral decay properties of the underlying physical operators. The parameter matrices in our dataset, generated via Gaussian Random Fields (GRF) or representing physical media, typically exhibit smoothness or piecewise smoothness corresponding to Sobolev regularity $H^{s}$ with $s>0$. Harmonic analysis dictates that the spectral energy of such functions decays polynomially or exponentially with frequency magnitude $|\mathbf{k}|$. Specifically, the Fourier coefficients satisfy $|\hat{P}_{\mathbf{k}}|\lesssim|\mathbf{k}|^{-s}$. Therefore, the truncation error is bounded by the tail energy of the spectrum:
$\epsilon_{ij}=\sum_{|\mathbf{k}|>p_{0}}|\Delta\hat{P}_{\mathbf{k}}|^{2}\approx\mathcal{O}(p_{0}^{-2s+d}),$
where $d$ is the spatial dimension. This decay rate implies that the high-frequency components contribute negligibly to the global topological distance between operators. This is also the fundamental reason why truncated FFT sorting is effective.